OpenStudy (anonymous):

How do I evaluate: 4^-1 + 2^-3?

OpenStudy (lgbasallote):

\[4^{-1} = \frac{1}{4}\] \[2^{-3} = \frac{1}{2^3} = \frac{1}{8}\]

OpenStudy (lgbasallote):

therefore \[4^{-1} + 2^{-3} = \frac{1}{4} + \frac{1}{8}\] got it?

mathslover (mathslover):

\[1/4 + 1/2^3\]\[1/4+1/8\]\[2+1/8\]\[3/8\]

OpenStudy (anonymous):

ohhh.. special thanks to @mathslover

mathslover (mathslover):

my pleasure